# Rollbaccine: Herd Immunity against Storage Rollback Attacks in TEEs (Technical Report)

### v1

@@ 7.4. Crash consistency and recovery
Figure 7. Recovery latency.
In this section, we simulate rollback attacks on both the primary and the backup in order to analyze the latency introduced by recovery and verify that Rollbaccine can always recover to a consistent state.
@@
We break the latency of recovery into three main phases in Figure 7: startup, hash transfer, and disk verification. Startup time (with “/” stripes) depends on whether Azure physically restarts the machine or redeploys it on a fresh VM; the decision is out of our control. In our experiment, the primary was physically restarted, and the backup was redeployed, taking 655 and 60 seconds respectively. Hash transfer (the thin gray line) is the time it takes for the recovering node to receive the 2.4GB in-memory integrity metadata from the other node; this takes 11 seconds in both tests. Disk verification (with “\” stripes) is the time it takes for the recovering node to read its the entire disk and perform integrity checks, recovering corrupted pages from the other node when necessary; this takes around 395 seconds in both tests, amounting to 1.5GB/s. This verification latency is unavoidable for any integrity-preserving application and is comparable to the 600 seconds it takes for dm-crypt + dm-integrity to format the disk.
We then test the correctness of Rollbaccine by simulating crashes and verifying the consistency of mounted file systems with ACE (crashMonkeyAndACE, ) and xfstests, standard tools for testing crash consistency. We generate and evaluate 577 tests on ext4 mounted over Rollbaccine. Rollbaccine passes all tests.
@@ 8. Related Work
Rollback resistance. Existing application-agnostic solutions for rollback resistance either sacrifice automatability or generality. Nimble (nimble, ) modifies applications to use its API in order to determine when data must be replicated. Narrator (narrator, ) assumes that applications are deterministic based on input ordering; accommodating non-deterministic applications requires recording executions for deterministic playback with high performance costs (dOS, ).
Rollback detection. Solutions that use hashes to verify integrity, but do not keep a backup of the data, are rollback detecting but not resistant (vpfs, ; mlsdisk, ; dm-x, ; lcm, ; rote, ).
Rollback resistance or detection has also been manually integrated into applications such as consensus protocols (engraft, ; ccf, ) and databases (speicher, ; treaty, ; fastver, ). SVR3 (svr3, ) introduces a variant of Raft (raft, ) that is safe even against physical attackers, assuming memory can only be rolled back at a page-level granularity.
Device mappers. Existing Linux device mappers offer some functionality to enforce confidentiality or integrity of disk. dm-crypt (dmCrypt, ) paired with dm-integrity (dmIntegrity, ) or dm-verity (dmVerity, ) can provide confidentiality and integrity in the presence of benign, random disk corruptions, but the integrity metadata on disk is vulnerable to rollbacks. drbd (drbd, ) replicates blocks but cannot detect attacks or compare the freshness of disks.
File system semantics. Prior work has explored substituting persistent file system operations for fault-tolerant replication (semanticAwareReplication, ) outside the context of rollback attacks. Assise (assise, ) uses this strategy for a NVM-backed network file system in order to reduce latency. SCFS (scfs, ) and drbd (drbd, ) allow users to toggle between replication schemes to replace disk persistence, while Gaios (paxosDisk, ) introduces replacements for file-related system calls that replicate to Paxos state machines. Blizzard (blizzard, ) replicates disk but acknowledges flushes before replication, breaking semantics in order to reduce latency.
@@ 9. Conclusion
Rollbaccine provides general, automatic, and low-overhead rollback resistance in a field where high performance and security are traditionally only achievable through careful code modifications. Rollbaccine achieves this by marrying the inherent asynchrony and concurrency of disk consistency with fault tolerant replication. Rollbaccine’s low overhead and generality leads us to believe that it can be transparently integrated into cloud storage systems with minimal effort.
@@ Acknowledgements
We thank Amaury Chamayou and Eddy Ashton for assistance with CCF (ccf, ) and AMD SEV-SNP, Vijay Chidambaram for pointing us to buffered durable linearizability (durableLinearizability, ), Alex Miller for double-checking persistence flag semantics, Srinath Setty for reviewing our Nimble (nimble, ) experiments, Shadaj Laddad and Tyler Hou for debugging memory reordering issues, Darya Kaviani for guiding us on cryptography, Ittai Abraham for suggesting we use $f+1$ instead of $12f+$ 1 nodes, Joe Hellerstein and Dimitra Giantsidi for additional feedback, and Mic Bowman, Michael Steiner, and Bruno Vavala for design discussions. This work was supported by gifts from AMD, Anyscale, Google, IBM, Intel, Microsoft, Mohamed Bin Zayed University of Artificial Intelligence, Samsung SDS, Uber, VMware, and Ripple.
@@ Appendix A Recovery
Recovery protocol. Our recovery protocol is based on the reconfiguration protocol from Matchmaker Paxos (matchmakerPaxos, ), with CCF (ccf, ) tracking configurations as the matchmakers.
@@
[…]
We optimize reconfiguration by omitting hashes from P1b and disk from both P1b and Reconfig, only requesting them when necessary. The recovering node first requests hashesd only from the designated node. It then performs an integrity scan over its local disk using hashesd, and only if any individual pages do not pass the integrity check, requests the page from the designated node. If the designated node fails during this process, the hashes are requested from another designated node (there must be another, since there are at most $f$ failures and each configuration has $f+1$ nodes), and the integrity scan is restarted with the new hashes. The recovering node then sends Reconfig to the other nodes in confi, which also perform integrity scans and request corrupted pages from the recovering node. If the designated node is also a node in confi, then it does not need to process Reconfig. This is the case for any recovery that replaces a single crashed node.
@@
Once recovery is complete, old configurations can be removed from allConf in CCF through a garbage collection protocol (matchmakerPaxos, ) and safely shut down.
@@ Appendix B Correctness
We first show that the backups respects any happens-before relationships on the primary. Writes are assigned writeIndex by the active primary based on invocation order. By definition of happens-before $V_{1}\prec V_{2}$ is only possible if $V_{1}$ precedes $V_{2}$, which implies that writeIndex of $V_{1}$ is also less than writeIndex of $V_{2}$. Therefore, if a backup submitted $V_{2}$ to disk, it must have already submitted $V_{1}$; formally, $V_{2}\in\mathcal{E}_{2}$ implies $V_{1}\in\mathcal{E}_{2}$.
We now show that the backups must contain all completed synchronous writes. The primary does not return synchronous writes to the application until the backups acknowledge that they have received that write and all prior writes with lower writeIndexes. Formally, $W_{res}(b,val,sync)\in\mathcal{E}_{1}$ implies $W_{res}(b,val,sync)\in\mathcal{E}_{2}$ if $sync$ contains REQ_FUA or REQ_PREFLUSH. By the definition of durable cut, $\mathcal{E}_{2}$ is a durable cut of $\mathcal{E}_{1}$, therefore $\mathcal{D}\mathcal{E}_{2}$ is a durable cut of $\mathcal{D}\mathcal{E}_{1}$. ∎
@@
Reconfiguration follows the protocol of Matchmaker Paxos (matchmakerPaxos, ). The proof can be derived from that of Matchmaker Paxos; we provide its intuition here.
@@
In the inductive case, if $\texttt{ballot}_{x}=\texttt{ballot}_{y}+i+1$, then there may be at most $i$ highest-ballot configuration in allConf. By the induction hypothesis, since there has been no active primaries since the configuration associated with balloty, no writes could have been made to disk, and each primary and backup must have synchronized their disks from either primary $y$, its backups, or some machine with state equivalent to those machines, and so must primary $x$. ∎

### v2

@@ 7.4. Crash Consistency and Recovery
Figure 8. Recovery latency. Each phase is labeled with its latency on top; hash transfer (unlabeled) takes 11 seconds in both experiments.
We simulate rollback attacks on both the primary and the backup in order to analyze recovery latency and correctness.
@@
We break the latency of recovery into three main phases in Figure 8: startup, hash transfer, and disk verification. Startup time depends on whether Azure physically restarts the machine or redeploys it on a fresh VM; the decision is out of our control. In our experiment, the primary was physically restarted, and the backup was redeployed, taking 655 and 60 seconds respectively. Hash transfer is the time it takes for the recovering node to receive the 2.4GB in-memory integrity metadata from the other node; this takes 11 seconds in both tests. Disk verification is the time it takes for the recovering node to read its the entire disk and perform integrity checks, recovering corrupted pages from the other node when necessary; this takes around 395 seconds in both tests, amounting to 1.5GB/s. This verification latency is unavoidable for any integrity-preserving application and is comparable to the 600 seconds it takes for dm-crypt + dm-integrity to format the disk. Recovery time increases linearly with the amount of corrupted disk; 100GB, 300GB, and 600GB of corrupted disk takes an additional 316, 951, and 1941 seconds respectively to recover, around 0.31GB/s.
We then test the correctness of Rollbaccine by simulating crashes and verifying the consistency of mounted file systems with ACE (crashMonkeyAndACE) and xfstests, standard tools for testing crash consistency. We generate and evaluate 577 tests on ext4 mounted over Rollbaccine. Rollbaccine passes all tests.
@@ 8. Related Work
Solutions against rollback attacks. No existing solution is simultaneously general, automatic, and rollback resistant.
Unlike Rollbaccine, which relies on device mappers for automation, Nimble (nimble) sacrifices automation, requiring applications to use its new API for rollback resistance. What it gains in return are (1) the ability to detect rollback attacks without cloud provider buy-in, and (2) a straightforward implementation over Intel SGX. Rollbaccine, in contrast, can (1) only protect against rollback attacks for systems mounted over its device mapper, and (2) requires a subset of the Linux kernel to be a part of the TCB (Trusted Computing Base), which is already part of the assumption for VM-based TEEs but requires significant reimplementation for Intel SGX (graphene; scone).
Narrator (narrator) sacrifices generality, requiring deterministic execution; accommodating non-determinism requires recording executions with high performance costs (dOS). The remaining solutions sacrifice resistance, using hashes or counters to verify integrity without keeping a backup of the data (vpfs; mlsdisk; dm-x; lcm; rote; crisp).
Device mappers. Existing Linux device mappers offer some functionality to enforce confidentiality or integrity of disk. dm-crypt (dmCrypt) paired with dm-integrity (dmIntegrity) or dm-verity (dmVerity) can provide confidentiality and integrity in the presence of benign, random disk corruptions, but the integrity metadata on disk is vulnerable to rollbacks.
File system semantics. Prior work has explored substituting persistent file system operations for fault-tolerant replication (semanticAwareReplication) outside the context of rollback attacks. Assise (assise) uses this strategy for a NVM-backed network file system in order to reduce latency. SCFS (scfs) and drbd (drbd) allow users to toggle between replication schemes to replace disk persistence, while Gaios (paxosDisk) introduces replacements for file-related system calls that replicate to Paxos state machines. Blizzard (blizzard) replicates disk but acknowledges flushes before replication, breaking semantics in order to reduce latency.
@@ 9. Conclusion
Rollbaccine provides general, automatic, low-overhead rollback resistance by marrying the inherent asynchrony and concurrency of disk consistency with fault tolerant replication. Rollbaccine’s low overhead and generality leads us to believe that it can be transparently integrated into cloud storage systems with minimal effort.
@@ Acknowledgements
This work would not be possible without support from Amaury Chamayou, Eddy Ashton, Joe Hellerstein, Ittai Abraham, Mic Bowman, Michael Steiner, Bruno Vavala, Vijay Chidambaram, Srinath Setty, Shadaj Laddad, Dimitra Giantsidi, Alex Miller, Tyler Hou, and Darya Kaviani. This work was supported by gifts from AMD, Anyscale, Google, IBM, Intel, Microsoft, Mohamed Bin Zayed University of Artificial Intelligence, Samsung SDS, Uber, VMware, and Ripple.
@@ Appendix A Recovery
Recovery protocol. Our recovery protocol is based on the reconfiguration protocol from Matchmaker Paxos (matchmakerPaxos), with CCF (ccf) tracking configurations as the matchmakers.
@@
After receiving at least 1 P1b from each configuration, the recovering node knows that no prior configuration can make progress and now selects the designated node to recover its state from. The designated node $d$ is the node with the highest $(\texttt{ballot},\texttt{writeIndex})$ pair, ordered lexicographically. The recovering node replaces its disk with diskd and sets its hashes to hashd. It then uses that hash and disk to update other nodes in its new configuration confi by sending Reconfig<seenBalloti, hashesd, diskd, writeIndexd>; those nodes replace their own hashes and disks similarly.
We optimize reconfiguration by omitting hashes from P1b and disk from both P1b and Reconfig, only requesting them when necessary. The recovering node first requests hashesd only from the designated node. It then performs an integrity scan over its local disk using hashesd, and only if any individual pages do not pass the integrity check, requests the page from the designated node. If the designated node fails during this process, the hashes are requested from another designated node (there must be another, since there are at most $f$ failures and each configuration has at least $f+1$ nodes), and the integrity scan is restarted with the new hashes. The recovering node then sends Reconfig to the other nodes in confi, which also perform integrity scans and request corrupted pages from the recovering node. If the designated node is also a node in confi, then it does not need to process Reconfig. This is the case for any recovery that replaces a single crashed node.
@@
Once recovery is complete, old configurations can be removed from allConf in CCF through a garbage collection protocol (matchmakerPaxos) and safely shut down.
@@ Appendix B Correctness
We first show that the backups respects any happens-before relationships on the primary. Writes are assigned writeIndex by the active primary based on invocation order. By definition of happens-before, $V_{1}\prec V_{2}$ is only possible if $V_{1}$ precedes $V_{2}$, which implies that writeIndex of $V_{1}$ is also less than writeIndex of $V_{2}$. Therefore, if a backup submitted $V_{2}$ to disk, it must have already submitted $V_{1}$; formally, $V_{2}\in\mathcal{E}_{2}$ implies $V_{1}\in\mathcal{E}_{2}$.
We now show that the backups must contain all completed synchronous writes. The primary does not return synchronous writes to the application until the backups acknowledge that they have received that write and all prior writes with lower writeIndexes. Formally, $W_{res}(b,val,sync)\in\mathcal{E}_{1}$ implies $W_{res}(b,val,sync)\in\mathcal{E}_{2}$ if syncsync contains REQ_FUA or REQ_PREFLUSH. By the definition of durable cut, $\mathcal{E}_{2}$ is a durable cut of $\mathcal{E}_{1}$, therefore $\mathcal{D}\mathcal{E}_{2}$ is a durable cut of $\mathcal{D}\mathcal{E}_{1}$. ∎
@@
Reconfiguration follows the protocol of Matchmaker Paxos (matchmakerPaxos). The proof can be derived from that of Matchmaker Paxos; we provide its intuition here.
@@
In the inductive case, $\texttt{ballot}_{x}=\texttt{ballot}_{y}+i+1$. Because there has been no active primaries since the configuration associated with balloty, no writes could have been made to disk, and each primary and backup must have synchronized their disks from either primary $y$, its backups, or some machine with state equivalent to those machines, and so must primary $x$. ∎